# LSMS: Language-guided Scale-aware MedSegmentor for Medical Image Referring Segmentation

### v2

@@ 3.3 Scale-aware Vision-Language Attention
As depicted in Figure 4(b), our proposed attention mechanism, termed Scale-aware Vision-Language Attention (SVLA), initially captures preliminary visual features through a convolution operation, and then employs Visual Knowledge Branch and the Language-Guided Branch to model rich visual knowledge and visual-linguistic relationships, followed by a $1\times 1$ convolution operation to learn the interplay between the branches. In $i$-th stage, with visual input $V_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ and language input $L\in\mathbb{R}^{C_{l}\times T}$, we obtain the preliminary visual feature map $V_{i}^{pre}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ by the formula $V_{i}^{pre}=Conv_{5\times 5}(V_{i})$.
@@
We employ the Language-Guided Branch to model relationships between linguistic information and various visual coordinates, facilitating the guidance of lesion localization in complex visual environment. The steps to obtain language-guided knowledge $Att_{i}^{L}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ are as follows:
@@
Table 1: Detailed settings of different stages in our LSMS.
Stage 1 Stage 2 Stage 3 Stage 4
number of blocks ($N_{i}$) 3 3 5 2
channel 64 128 320 512
@@
3.4 Full-Scale Decoder
In order to integrate linguistic cues for comprehending complex medical visual environment, a comprehensive cross-scale understanding of visual-linguistic contexts is necessary. Therefore, we devised a Full-Scale Decoder (FSD) to capture high-level semantics following the encoder. Unlike previous methods [41, 37, 10] with sequential structures, we globally process multi-modal features of various scales after aligning their positions. We employ Position Alignment to map them to the same feature map size of $F_{1}$ while retaining their original channel numbers. Subsequently, we concatenate features from different scales, pass them through a lightweight InterScale block for globally modeling multi-scale contexts, and finally generate segmentation predictions. The process is as follows:
$S_{1},S_{2},S_{3},S_{4}=PositionAlign(F_{1},F_{2},F_{3},F_{4}),$ (11)
$Out=Seg(InterScale(Concat[S_{2},S_{3},S_{4}])),$ (12)
@@ 3.4 Full-Scale Decoder
It is noteworthy that we exclusively utilize features generated from $2,3,4$-th stages for global decoding. This choice is informed by the observation that the shallow features from the first stage exhibit a lower degree of visual-linguistic consistency, encompassing redundant information from irrelevant regions in the images. This redundancy hampers lesion localization and hinders segmentation. The efficacy of this design will be validated in the Ablation Study (Section 4.5) and further elucidated through visualization analysis (Section 4.6) to comprehend the disparities and characteristics of features from different stages.
4 Experiments
4.1 Datasets
RefHL-Seg represents the first dataset tailored specifically for the MIRS task. RefHL-Seg comprises 2,283 abdominal CT slices from 231 cases, predominantly featuring lesions such as Hemangiomas, Liver Cysts, Hepatocellular Carcinomas (HCC), Focal Nodular Hyperplasia, and Metastasis. With collaboration from radiology experts, we meticulously annotated all liver lesions for the first time, providing detailed descriptions of their locations and morphologies. Each lesion’s textual annotation includes information such as liver segment (location), diameter, shape, boundary, enhancement pattern, lesion type, and more. An exemplary textual annotation containing all relevant information is provided as follows: "A vascular tumor in the VI segment of the liver, with a diameter of 7.5mm, irregular shape, clear boundary, and no ring enhancement." Additionally, experiments will involve testing language expressions that contain only partial information sufficient for lesion localization, such as "A vascular tumor in the VI segment of the liver, with an irregular shape."
The MosMedData+ [23, 14] and QaTa-COV19 [5] datasets are established datasets for conventional medical image segmentation tasks. [20] extended the textual annotations of these datasets, serving for the evaluation of Medical Image Segmentation with Textual Prompts. The MosMedData+ dataset comprises 2,729 CT scan slices of lung infections, primarily containing information about the location of lung infections and the number of infected regions. The QaTa-COV19 dataset consists of 9,258 chest X-ray images manually annotated with COVID-19 lesions, focusing on whether both lungs are infected, the quantity of lesions, and the approximate location of the infected regions.
4.2 Evaluation Metric
Table 2: Experimental results on the RefHL-Seg, QaTa-COV19 and MosMedData+ datasets in terms of Dice and mIoU. The best scores are in red, and the secondbest scores are in blue.
@@ 4.2 Evaluation Metric
U-Net [29] CNN $\times$ 14.8 50.3 48.67 37.89 79.02 69.46 64.60 50.73
UNet++ [43] CNN $\times$ 74.5 94.6 51.84 40.27 79.62 70.25 71.75 58.39
AttUNet [25] CNN $\times$ 34.9 101.9 50.67 40.11 79.31 70.04 66.34 52.82
nnUNet [16] CNN $\times$ 19.1 412.7 51.63 41.89 80.42 70.81 72.59 60.36
TransUNet [3] Hybrid $\times$ 105 56.7 52.33 44.76 78.63 69.13 71.24 58.44
Swin-Unet [2] Hybrid $\times$ 82.3 67.3 51.34 44.26 78.07 68.34 63.29 50.19
UCTransNet [33] Hybrid $\times$ 65.6 63.2 54.87 44.38 79.15 69.60 65.90 52.69
LViT [20] Hybrid $\times$ 28.0 54.0 54.84 44.65 81.12 71.37 72.58 60.40
LSMS Hybrid $\times$ 8.78 8.91 58.75 49.39 82.14 72.07 73.14 61.76
@@
SLViT [26] Hybrid 133.6 52.2 71.94 62.83 80.68 71.57 73.08 60.11
@@
LSMS (1/4) Hybrid 25% 8.78 8.91 77.57 68.95 83.39 75.09 76.47 65.05
LSMS (1/2) Hybrid 50% 8.78 8.91 77.63 69.14 83.44 75.13 76.98 65.71
LSMS Hybrid 100% 8.78 8.91 78.14 70.22 84.05 75.70 77.88 66.83
Table 3: Ablation studies on the RefHL-Seg val set. The optimal scores are highlighted in red.
@@
(a) Kernel size of a single ConvUnit in SVLA
@@
76.34 66.96
76.52 67.03
76.44 67.17
77.10 68.88
78.14 70.22
Dice mIoU
(c) Ablation on FSD
LSMS (w/o FSD) 74.86 65.64
LSMS (w/ FSD) 78.14 70.22
@@
$S_{4}$-Head-MLP 75.82 67.39
$S_{1},S_{2},S_{3},S_{4}$-MLP-Concat-MLP 77.07 68.48
$S_{1},S_{2},S_{3},S_{4}$-Concat-Head-MLP 77.95 69.33
$S_{1},S_{2},S_{3},S_{4}$-MLP-Concat-Head-MLP 76.35 68.21
@@
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$
77.82 66.39
78.71 67.26
77.83 66.92
78.14 70.22
77.95 69.33
4.3 Implementation Details
The proposed LSMS is implemented using PyTorch, leveraging the BERT implementation from the HuggingFace Transformer library [36]. Regarding dataset partitioning, we separated the original training set into training and validation sets. For the convolutional layers in the Visual Knowledge Branch of SVLA, we initialized the weights using SegNeXt [12] pre-trained on ImageNet-22K [13]. The language encoder of our model was initialized with the official pre-trained BERT weights, consisting of 12 layers with a hidden size of 768. The number of encoder blocks and feature dimensions for each stage are presented in Table 1. For the convolutional branch of SVLA, we used kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$. The remaining weights in our model were randomly initialized.
@@ 4.3 Implementation Details
4.4 Comparison with the State-of-the-Arts
Classical Medical Image Segmentation
@@ 4.4 Comparison with the State-of-the-Arts
We compared the performance of LSMS with existing segmentation models on the RefHL-Seg, QaTa-COV19, and MosMedData+ datasets, as shown in Table 2. In scenarios where language input is not utilized, we observed that LSMS outperformed all existing methods in classical medical image segmentation task while minimizing computational costs. This underscores the efficiency and superiority of LSMS in understanding medical visual environment.
Medical Image Segmentation with Textual Prompts
@@
The QaTa-COV19 and MosMedData+ datasets are primarily utilized for training and validating Medical Image Segmentation with Textual Prompts. In Table 2, our LSMS achieves state-of-the-art (SOTA) performance for this task, particularly improving Dice by 3.21% and mIoU by 5.50% on the MosMedData+ dataset. This indicates that LSMS, with the assistance of textual descriptions, can complement visual information with linguistic cues, leading to more accurate segmentation of all lesions in the images and thereby achieving superior outcomes.
Medical Image Referring Segmentation
@@
The RefHL-Seg dataset we constructed serves the specific purpose of training and validating the MIRS task. Each sample in the dataset necessitates the model to segment the particular lesion indicated by the reference expression, thus rendering the performance of all methods suboptimal when text input is absent. When provided with complete inputs containing both text and images, our LSMS significantly improves accuracy compared to all existing models. Specifically, compared to the previous SOTA method LViT [20], LSMS achieves an increase of 6.66% in Dice and 7.80% in mIoU. This improvement stems from LSMS’s enhanced understanding of the complex medical visual environment and its precise localization capability based on linguistic cues. In Table 2, LSMS (1/4) and LSMS (1/2) denote the models tested with only 25% and 50% of the textual input, respectively, by omitting portions such as size or shape descriptions. The results indicate that LSMS still outperforms existing methods significantly. This underscores that our LSMS, trained to precisely locate specified lesions in images with minimal textual assistance, exhibits outstanding language-guided lesion localization capability.
@@
Figure 5: Visualization of the feature maps from different stages in LSMS. The red regions denote the Ground Truth, while the green regions represent the segmentation results of our LSMS.
4.5 Ablation Study
Kernel size of a single ConvUnit in SVLA
@@ 4.5 Ablation Study
We have meticulously evaluated the performance of using different convolutional kernels in SVLA on the validation set of RefHL-Seg, as illustrated in Table 3(a). $d_{j}=N$ indicates the ConvUnit containing a $1\times N$ convolution and a $N\times 1$ convolution. The strip-like convolution kernels aim to obtain detailed local visual information with low costs. We employed a single ConvUnit in SVLA to evaluate the impact of various convolution kernel sizes $d_{j}$. In Table 3(a), sizes 7 and 21 show superior performance among those with comparable computational costs. The performance of medium-size convolution kernels is similar, so we choose size 11 due to its lower computational cost. The utilization of diverse convolution kernels can aid in capturing features from varying receptive fields, which helps in extracting rich local visual features. We opted for kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$ as alternatives.
@@
To assess the effectiveness of the FSD, we compared the performance of the complete LSMS with LSMS (w/o FSD), and the results are reported in Table 3(c). Experimental findings indicate that removing the FSD module leads to performance degradation, with Dice decrease of 3.28% and mIoU decrease of 4.58%, respectively. Furthermore, FSD incurs only a modest increase of 2.1M parameters and 1.7G Flops, yet significantly enhances performance. This underscores its efficiency in boosting performance by globally modeling multi-modal information.
@@
To validate the effectiveness of the individual components in the design of FSD, we conducted ablation experiments on its constituents, as presented in Table 3(d). It is evident that relying solely on features from the final stage is insufficient, and while adding MLP layers before concatenation increases complexity, it results in the loss of valuable information from each stage. The incorporation of the Hamburger Head design enhances the ability to globally model multi-scale information, consequently improving segmentation performance. The $S_{1},S_{2},S_{3},S_{4}$-Concat-Head-MLP design demonstrates the best performance.
@@
Given the multi-modal features from different stages after Position Alignment, FSD concatenates them and performs joint refinement in a single forward pass. Here, $S_{i}$, $i\in\{1,2,3,4\}$, denotes the multi-modal features input to FSD from the $i$-th stage. Table 3(e) compares multiple input sequences, confirming the value of multi-scale interaction for global reasoning. As shown, $S_{2},S_{3},S_{4}$ exhibit optimal performance for FSD. This superiority is attributed to the insufficient depth of interaction between visual information and linguistic cues in the shallow feature $S_{1}$, which include irrelevant information and hindering the localization and segmentation of specified lesions. Conversely, the latter three layers demonstrate strong visual-linguistic consistency, facilitating favorable predictions. Visualization analysis of each feature map is detailed in Section 4.6.
4.6 Visualization Analysis

### v3

@@ III-C Scale-aware Vision-Language Attention
As depicted in Fig. 5(b), our proposed attention mechanism, termed Scale-aware Vision-Language Attention (SVLA), initially captures preliminary visual features through a convolution operation, and then employs Visual Knowledge Branch and Language-Guided Branch to model rich visual knowledge and visual-linguistic relationships, followed by a $1\times 1$ convolution operation to learn the interplay between the branches. In $i$-th stage, with visual input $V_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ and language input $L\in\mathbb{R}^{C_{l}\times T}$, we obtain the preliminary visual feature map $V_{i}^{pre}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ by the formula $V_{i}^{pre}=Conv_{5\times 5}(V_{i})$.
@@
Figure 5: (a) An illustration of the encoder block in the Language-Guided Vision Encoder. (b) An illustration of the Scale-aware Vision-Language Attention module.
@@
We employ the Language-Guided Branch to model relationships between linguistic information and various visual coordinates, facilitating the guidance of lesion localization in the complex visual environment. The steps to obtain language-guided knowledge $Att_{i}^{L}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ are as follows:
@@
III-D Full-Scale Decoder
To accommodate the complexity of medical visual environment, a comprehensive cross-scale understanding of visual-linguistic contexts is necessary. Therefore, we devised a Full-Scale Decoder (FSD) to capture high-level semantics following the encoder. Unlike previous methods [29, 30, 31] with sequential structures, we globally process multi-modal features of various scales after aligning their positions. We employ Position Alignment to map them to the same feature map size of $F_{1}$ while retaining their original channel numbers. Subsequently, we concatenate features from different scales, pass them through a lightweight InterScale block for globally modeling multi-scale contexts, and finally generate segmentation predictions. The process is as follows:
$P_{1},P_{2},P_{3},P_{4}=PositionAlign(F_{1},F_{2},F_{3},F_{4}),$ (11)
$Out=Seg(InterScale(Concat[P_{2},P_{3},P_{4}])),$ (12)
@@ III-D Full-Scale Decoder
It is noteworthy that we exclusively utilize features generated from $2,3,4$-th stages for global decoding. This choice is informed by the observation that the shallow features from the first stage exhibit a lower degree of visual-linguistic consistency, encompassing redundant information from irrelevant regions in the images. This redundancy hampers lesion localization guided by linguistic cues and interferes with the segmentation of target lesions in the visual environment. The efficacy of this design will be validated in the Ablation Study (Sec. IV.E). Further visualization analysis (Sec. IV.F) will be conducted to explore the disparities and characteristics of features from different stages.
III-E Loss Function
We design a specialized loss function to constrain the model’s training, which includes the Segmentation Loss $\mathcal{L}_{Seg}$ for the segmentation results and the Vision-Language Contrastive Loss $\mathcal{L}_{Con}$ for the linguistic features and final multi-modal features.
Specifically, we perform dilation and erosion operations on the ground truth label $M^{gt}$ to obtain the boundary region $E$, defined as $E=dilate(M^{gt})-erode(M^{gt})$. Here, $dilate(\cdot)$ expands and $erode(\cdot)$ contracts the object’s boundaries. A weight $\lambda$ is then assigned to the edge region where $E^{i}=1$. The loss function for the segmentation mask is defined as:
$\mathcal{L}_{Seg}=\mathcal{L}_{focal}(M,M^{gt})+W\odot\mathcal{L}_{dice}(M,M^{gt}),$ (13)
where $W^{i}=\lambda E^{i}+(1-E^{i})$, $\mathcal{L}_{focal}$ and $\mathcal{L}_{dice}$ are focal loss [33] and dice loss [34].
Furthermore, $\mathcal{L}_{Con}$ aligns the visual knowledge of the target region with the linguistic information while simultaneously repelling the visual representations of the target and non-target regions, as illustrated in Figure 3(b). In the figure, $S_{align}$ measures the similarity between the features of positive pixels and the average linguistic feature, while $S_{repel}$ quantifies the dissimilarity between the features of positive and negative pixels. The formula for $\mathcal{L}_{Con}$ is as follows:
$\mathcal{L}_{Con}$ $=-\frac{1}{|\mathcal{P}|}\sum_{i}^{|\mathcal{P}|}\frac{e^{(p_{i}\cdot L_{avg}/\tau)}}{e^{(p_{i}\cdot L_{avg}/\tau)}+\sum_{j}^{|\mathcal{N}|}e^{(p_{i}\cdot n_{j}/\tau)}},$ (14)
where $p_{i}$ represents the feature of the $i$-th positive pixel in the positive pixel set $\mathcal{P}$, $n_{j}$ denotes the feature of the $j$-th negative pixel in the negative pixel set $\mathcal{N}$, and $\tau$ is a hyperparameter that controls the sharpness of the probability distribution. $L_{avg}$ denotes the average pooled linguistic feature, computed as $L_{avg}={proj}\left(\frac{1}{T}\sum^{T}L_{t}\right)$. $L_{t}$ is the $t$-th linguistic token.
Our joint training loss is defined as follows:
$\mathcal{L}=\alpha\mathcal{L}_{Con}+\mathcal{L}_{Seg},$ (15)
where $\alpha$ is a hyperparameter.
TABLE II: Experimental results on the RefHL-Seg, QaTa-COV19 and MosMedData+ datasets in terms of Dice and mIoU. The best scores are in red, and the secondbest scores are in blue.
@@ III-E Loss Function
U-Net [1] CNN 14.8 50.3 48.67 37.89 79.02 69.46 64.60 50.73
UNet++ [2] CNN 74.5 94.6 51.84 40.27 79.62 70.25 71.75 58.39
AttUNet [35] CNN 34.9 101.9 50.67 40.11 79.31 70.04 66.34 52.82
nnUNet [36] CNN 19.1 412.7 51.63 41.89 80.42 70.81 72.59 60.36
TransUNet [3] Hybrid 105 56.7 52.33 44.76 78.63 69.13 71.24 58.44
Swin-Unet [7] Hybrid 82.3 67.3 51.34 44.26 78.07 68.34 63.29 50.19
UCTransNet [4] Hybrid 65.6 63.2 54.87 44.38 79.15 69.60 65.90 52.69
LViT [5] Hybrid 28.0 54.0 54.84 44.65 81.12 71.37 72.58 60.40
LSMS Hybrid 8.78 8.91 58.75 49.39 82.14 72.07 73.14 61.76
@@
SLViT [10] Hybrid 103.4 50.1 71.94 62.83 80.68 71.57 73.08 60.11
Prompt-RIS [27] Hybrid 225.9 108.7 72.11 63.08 81.13 71.94 74.12 60.97
@@
RecLMIS [8] CNN 23.7 24.1 73.01 63.59 85.22 77.00 77.48 65.07
LSMS (1/4) Hybrid 25% 8.78 8.91 77.63 69.02 84.97 76.98 77.46 66.02
LSMS (1/2) Hybrid 50% 8.78 8.91 78.33 69.76 85.21 77.08 77.65 66.24
LSMS Hybrid 100% 8.78 8.91 78.24 70.31 85.57 77.60 78.32 67.41
IV Experiments
IV-A Datasets
RLS, as a new task, lacks suitable benchmarks. The benchmarks established for Natural Image Referring Segmentation cannot be directly transferred to the medical domain due to fundamental differences in acquisition conditions and visual properties. Medical images differ from natural images in texture, contrast, and annotation complexity, requiring expert knowledge. Their targets often have unclear boundaries and varying sizes, necessitating multi-scale processing. To tackle these issues, we developed the Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset, specifically designed for the RLS task.
As the first dataset tailored specifically for RLS, RefHL-Seg comprises 2,283 abdominal CT slices from 231 cases, predominantly featuring lesions such as Hemangiomas, Liver Cysts, Hepatocellular Carcinomas (HCC), Focal Nodular Hyperplasia, and Metastasis. With collaboration from radiology experts, we meticulously annotated all liver lesions for the first time, providing detailed descriptions of their locations and morphologies. Each lesion’s textual annotation includes information such as liver segment (location), diameter, shape, boundary, enhancement pattern, lesion type, and more. Corresponding segmentation masks were delineated. An exemplary textual annotation containing all relevant information is provided as follows: “A vascular tumor in segment VI of the liver, with a diameter of 7.5mm, irregular shape, clear boundary, and non-ring enhancement.” Additionally, experiments will involve testing language expressions that contain only partial information sufficient for lesion localization, such as ”A vascular tumor in segment VI of the liver, with an irregular shape.”
The MosMedData+ [42, 43] and QaTa-COV19 [44] datasets are established for Classical Lesion Segmentation tasks. Recent study [5] extended the textual annotations of these datasets, serving for the evaluation of Text-Augmented Lesion Segmentation. The MosMedData+ dataset comprises 2,729 CT scan slices of lung infections, primarily containing information about the location of lung infections and the number of infected regions. The QaTa-COV19 dataset consists of 9,258 chest X-ray images manually annotated with COVID-19 lesions, focusing on whether both lungs are infected, the quantity of lesions, and the approximate location of the infected regions.
IV-B Evaluation Metric
TABLE III: Ablation studies on the RefHL-Seg val set. The optimal scores are highlighted in red.
@@ IV-B Evaluation Metric
(a) Kernel size of the single ConvUnit in SVLA
@@
76.75 67.34
76.89 67.45
76.65 67.22
77.23 68.91
78.34 70.31
(c) Effectiveness of FSD
LSMS (w/o FSD) 74.92 65.77
LSMS (w/ FSD) 78.34 70.31
@@
$P_{4}$-Head-MLP 76.03 67.58
$P_{1},P_{2},P_{3},P_{4}$-MLP-Concat-MLP 77.41 68.79
$P_{1},P_{2},P_{3},P_{4}$-Concat-Head-MLP 77.97 69.38
$P_{1},P_{2},P_{3},P_{4}$-MLP-Concat-Head-MLP 76.40 68.29
@@
$P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$
77.94 66.59
78.80 67.54
77.89 67.36
78.34 70.31
77.97 69.38
IV-C Implementation Details
The proposed LSMS is implemented using PyTorch, leveraging the BERT implementation from the HuggingFace Transformer library [45]. Regarding dataset partitioning, we separated the original training set into training and validation sets. For the convolutional layers in the Visual Knowledge Branch of SVLA, we initialized the weights using SegNeXt [46] pre-trained on ImageNet-22K [47]. The language encoder of LSMS was initialized with the official pre-trained BERT weights, consisting of 12 layers with a hidden size of 768. The number of encoder blocks and feature dimensions for each stage are presented in Tab. I. For the convolutional branch of SVLA, we used kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$. We set the default values for key hyperparameters as follows: $\alpha=0.125$, $\lambda=1.2$ and $\tau=0.1$. The remaining weights in our model were randomly initialized.
@@ IV-C Implementation Details
IV-D Comparison with the State-of-the-Arts
Classical Lesion Segmentation
@@ IV-D Comparison with the State-of-the-Arts
We compared the performance of LSMS with existing segmentation models on the RefHL-Seg, QaTa-COV19, and MosMedData+ datasets, as shown in Tab. II. In scenarios where Text is not utilized, we remove the language-related components from LSMS, such as the Language-Guided Branch in SVLA and the $\mathcal{L}_{Con}$ term in the loss function. We observed that LSMS outperformed all existing methods in Classical Lesion Segmentation task while minimizing computational costs. This underscores the efficiency and superiority of LSMS in understanding medical visual environment.
Text-Augmented Lesion Segmentation
@@
The QaTa-COV19 and MosMedData+ datasets have been extended with textual annotations for training and validating Text-Augmented Lesion Segmentation. In Tab. II, our LSMS achieves state-of-the-art (SOTA) performance. As the proportion of textual input increases, LSMS consistently improves across all evaluation metrics. This indicates that LSMS, with textual assistance, can complement visual information with linguistic information, leading to a more accurate segmentation of the lesions.
Referring Lesion Segmentation
@@
The RefHL-Seg dataset we constructed serves the specific purpose of training and validating the RLS task. Each sample in the dataset necessitates the model to segment the particular lesion indicated by the reference expression, thus rendering the performance of all methods suboptimal when text input is absent. Our LSMS serves as a novel baseline for the RLS task, demonstrating its significant advantages over methods applied to previous related tasks. To provide a comprehensive comparison, we evaluate LSMS against existing vision-language models on the newly proposed task and dataset. When provided with complete inputs containing both text and images, our LSMS significantly improves accuracy compared to all existing models. Specifically, LSMS achieves an increase of 8.29% over LViT [5] and 6.72% over RecLMIS [8] in mIoU, highlighting its effectiveness and superior performance in this new benchmark. This improvement stems from LSMS’s enhanced understanding of the complex medical visual environment and its precise localization capability based on linguistic cues. In Tab. II, LSMS (1/4) and LSMS (1/2) denote the models tested with only 25% and 50% of the textual input, respectively, by omitting portions such as size or shape descriptions. The results indicate that LSMS still outperforms existing methods significantly. This underscores that our LSMS, trained to precisely locate specified lesions in images with minimal textual guidance, exhibits outstanding language-guided lesion localization capability.
@@
Figure 6: Visualization of the feature maps from different stages in LSMS. The red regions denote the Ground Truth, while the green regions represent the segmentation results of our LSMS. In sample (b), for ease of observation, the key region within the image have been enlarged, with a white rectangular box serving as a reference for location.
IV-E Ablation Study
Kernel size of the single ConvUnit in SVLA
@@ IV-E Ablation Study
We have meticulously evaluated the performance of using different convolutional kernels in SVLA on the validation set of RefHL-Seg dataset, as illustrated in Tab. III(a). $d_{j}=N$ indicates the ConvUnit containing a $1\times N$ convolution and a $N\times 1$ convolution. The strip-like convolution kernels aim to obtain detailed local visual information with low costs. We employed a single ConvUnit in SVLA to evaluate the impact of various convolution kernel sizes $d_{j}$. In Tab. III(a), sizes 7 and 21 show superior performance among those with comparable computational costs. The performance of medium-size convolution kernels is similar, so we choose size 11 due to its lower computational cost. The utilization of diverse convolution kernels can aid in capturing features from varying receptive fields, which helps in extracting rich local visual features. Therefore, we selected kernel sizes of $d_{1}=7$, $d_{2}=11$, and $d_{3}=21$ as the default settings.
@@
To assess the effectiveness of the FSD, we compared the performance of the complete LSMS with LSMS (w/o FSD), and the results are reported in Tab. III(c). Experimental findings indicate that removing the FSD module leads to performance degradation, with Dice decrease of 3.42% and mIoU decrease of 4.54%, respectively. Furthermore, FSD incurs only a modest increase of 2.1M parameters and 1.7G Flops, yet significantly enhances performance. This underscores its efficiency in boosting performance by globally modeling multi-modal information.
@@
To validate the effectiveness of the individual components in the design of FSD, we conducted ablation experiments on its constituents, as presented in Tab. III(d). It is evident that relying solely on features from the final stage is insufficient, and while adding MLP layers before concatenation increases complexity, it results in the loss of valuable information from each stage. The incorporation of the Hamburger Head design enhances the ability to globally model multi-scale information, consequently improving segmentation performance. The $P_{1},P_{2},P_{3},P_{4}$-Concat-Head-MLP design demonstrates the best performance.
@@
Given the multi-modal features from different stages after Position Alignment, FSD concatenates them and performs joint refinement in a single forward pass. Here, $P_{i}$, $i\in\{1,2,3,4\}$, denotes the multi-modal features input to FSD from the $i$-th stage. Tab. III(e) compares multiple input sequences, confirming the value of multi-scale interaction for global reasoning. As shown, $P_{2},P_{3},P_{4}$ exhibit optimal performance for FSD. This superiority is attributed to the insufficient depth of interaction between visual information and linguistic cues in the shallow feature $P_{1}$, which include irrelevant information and hindering the localization and segmentation of specified lesions. Conversely, the latter three layers demonstrate strong visual-linguistic consistency, facilitating favorable predictions. Visualization analysis of each feature map is detailed in Sec. IV.F.
IV-F Visualization Analysis